# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

### v4

@@ 3. Methods: Interview Study
To hear firsthand from Reddit moderators and to deeply explore a new area of research, we chose to conduct qualitative research via semi-structured interviews. We designed an initial interview protocol based on our research questions and updated it as themes emerged in early interviews. Participants were allowed to choose a medium that was comfortable to them for the interview: video call, audio call, email exchange, or message exchange via the Reddit platform. Regardless of channel, we posed our research questions one at a time and followed up in a semi-structured way. Pseudonymity is an important feature of the Reddit platform, so we did not ask for or record participants’ real names or offline identifiers. This study was IRB-approved. All participants read and accepted a consent agreement permitting their responses to be reproduced in research reports, along with the participants’ associated subreddits. To ensure that we reliably represented participants’ views, we contacted participants for feedback and approval on their included quotes before publication, following CSCW community recommendations (mcdonald19). We were able to solicit, receive, and integrate all participants’ feedback.
3.1. Crawl of Public Subreddit Rules
Our goal was to speak to moderators of communities who were actively governing the use of AIGC. To identify such communities, we launched a data collection effort to find subreddits with public community rules pertaining to AI use. We performed an automated crawl of the Reddit website in the summer of 2023 and extracted the text-based rules from all subreddits indexed by Reddit’s “Top Communities” list [^1] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We crawled $n=337,399$ subreddits and found $n=87,596$ with non-empty rules and English as a primary language. We then performed a text search on the extracted rules to identify subreddits with rules related to AIGC. We acknowledge that this method is imprecise, but we found it sufficient for our exploratory analysis.
We found that 4% of all subreddits in the Top Communities index had rules about AI, but these rules were much more common in larger subreddits: when we restricted our analysis to top 1% of subreddits according to subscriber count, 20% had rules governing the use of AI. We thus decided to build our sample starting with the largest subreddits as these appear to be the most affected by AI at this moment. We also heard about moderating smaller communities from interviewees who moderated several communities of different sizes.
@@ 3.2. Recruitment
We recruited and interviewed participants during the summer of 2023. We manually examined the rules of subreddits in descending order of subscriber count. When we found a subreddit with either an explicit rule governing the use of AI or a rule that might implicitly cover AI use we added it to our sample, avoiding communities who banned research outreach. In addition, we also looked for public moderator discussions concerning AIGC in other forums, such as the r/ModSupport subreddit, and added those moderators’ subreddits to our sample. We then messaged the moderators of each subreddit in our sample via Reddit’s Modmail [^2] https://support.reddithelp.com/hc/en-us/articles/210896606-What-is-Modmail- feature, sending Modmail recruitment messages to 60 subreddits. We used two additional techniques to recruit participants beyond the largest subreddits. First, we conducted snowball sampling by asking moderators to forward our recruitment message to other potential participants. Second, we posted our recruitment call on several social media platforms. To ensure respondents were actual moderators, we asked for their Reddit username and the subreddit that they moderate, then verified that the users were indeed listed as moderators via their Reddit user profile. Participants were offered \$20 gift-cards as compensation. Overall, we interviewed fifteen participants who collectively moderated over 100 different subreddits with sizes ranging from less than ten members to more than 32 million (see participant information in Table 1).
@@ 3.3. Interviews
The first author conducted semi-structured 60 minute interviews with moderators about their experience with AIGC. The interviews used these key guiding questions:
@@
    What concerns do moderators have about AIGC and how severe are they compared to their other community concerns?
@@
Interview transcripts were generated automatically from recordings and were manually reviewed and corrected by the first author. The first two authors then collaborated to iteratively code the interview content into categories, using an inductive thematic analysis method akin to Braun and Clarke’s Reflexive Thematic Analysis (braun21; Braun19). The two authors began by discussing themes that the first author observed while conducting the interviews and correcting the transcripts. From this conversation, an initial high-level set of codes emerged. Next, the first author open-coded all of the interview transcripts at the sentence-level, applying the initial codes and creating new ones where existing codes were insufficient. The two authors then reviewed the set of codes and refined them by merging similar codes and nesting those with a clear hierarchy. This round of coding produced no completely new themes, only new sub-codes, so the two authors agreed that the codes seemed fairly stable. Both authors then independently applied these codes to the same three interview transcripts and compared their results. Neither had created any new codes and the few cases of disagreement were resolved by further merging similar codes. This process yielded approximately 400 tagged phrases across all transcripts, which were grouped into the four high-level themes that we discuss below.
@@
1 r/todayilearned (32M) 25 Audio Call Amplifying Attacks High
2 r/food (23M) 25 Reddit Message Lowering Quality: More Off-Topic High
3 r/explainlikeimfive (23M) 3 Email Lowering Quality: Less Accurate Information High
4 r/explainlikeimfive (22M) 2 Audio Call Lowering Quality: Less Accurate Information High
5 r/WritingPrompts (17M) 3 Video Call Lowering Quality: More Off-Topic High
6 r/politics (8.3M) 25 Reddit Message Straining Relationships Medium
7 r/CryptoCurrency (6.7M) 8 Reddit Message Amplifying Attacks Medium
8 r/itookapicture (5.1M) 2 Email Lowering Quality: More Off-Topic Medium
9 r/NintendoSwitch (5.0M) 23 Email Lowering Quality: More Off-Topic Medium
10 r/Fantasy (3.4M) 1 Email Lowering Quality: More Off-Topic Medium
11 r/changemyview (3.4M) 2 Video Call Lowering Quality: Less Human Connection Low
12 r/changemyview (3.4M) 2 Audio Call Lowering Quality: Less Human Connection Low
13 r/AskHistorians (1.8M) 4 Video Call Lowering Quality: Less Accurate Information Medium
14 r/AskHistorians (1.8M) 1 Reddit Message Lowering Quality: Less Accurate Information Medium
15 r/GCTrading (22.7K) 3 Audio Call Amplifying Attacks Low
@@ 4. Findings
We share findings about subreddit moderators’ perspectives towards AIGC based on their personal moderation experiences, which center on four key themes: their general attitudes towards AIGC, the specific threats that they perceive AIGC poses to their communities, the ways that their communities are responding to the presence of AIGC, and the challenges of enforcing rules about AIGC.
4.1. Findings: General Attitudes Towards AIGC (RQ1)
While our participants acknowledged some potential benefits of AIGC, they reported generally opposing its use in their communities. Below we share insights into our participants’ various definitions of AIGC, then follow with details about their opposition to its use in their communities, and conclude this section with some potential benefits of AIGC that our participants also acknowledged.
4.1.1. Definitions of AIGC
Our interviews did not reveal a single consensus about the definition of “AIGC”, but we heard our participants use the term to refer to three different categories of content: images produced using a generative AI tool, text produced either entirely or partially with a generative AI tool, and posts made by automated accounts, or bots. Communities sometimes banned one of these categories of content while still allowing others.
We also heard different perspectives about what AIGC was problematic. Some participants categorically rejected AIGC, like a moderator of r/WritingPrompts [^3] To provide additional context, we introduce quotes with the name of the subreddit that the participant moderates. In the case where the participant moderates several subreddits, we include either the subreddit that the quote is referring to or, for more general quotes, the largest subreddit that the participant moderates. , who countered a community member’s argument that, “ChatGPT is another writing tool that authors are using and banning ChatGPT will be the same as banning a spellchecker,” with this response: “that argument is not good, because the level of effort required is clearly shifting quite dramatically.” Other participants acknowledged that AI-use can fall along a spectrum, from being used for minor editing to being used to completely write a post, and thus specific instances of AIGC are evaluated based on subjective characteristics. One participant who moderates r/itookapicture mentioned the importance of their perception of a poster’s [^4] We use the term poster to refer to both those posting top-level subreddit submissions as well as those posting comments in response to submissions. motive: “I don’t mind people using AI to better state their thoughts nor create art as long as it is not being used deceptively.” We also heard from participants who judged AIGC based on whether or not they deemed it to be the poster’s original work. For some participants, this was a matter of the degree to which AIGC still contained the poster’s own voice, while to others it was about AIGC being a poster’s “own intellectual contribution or content”:
@@ 4.1.1. Definitions of AIGC
The perceived intent and degree to which AIGC represents a poster’s own voice are subjective characteristics which could easily lead to disagreement about whether or not a specific example of AIGC is acceptable, which we revisit in later sections.
4.1.2. Opposition To AIGC
We dug deeper into the reasons behind our participants’ different notions of what makes AIGC problematic and found that our participants opposed AIGC in their communities for either ideological or practical reasons. Our participants voiced three different levels of objection to AIGC, which ranged from general to specific. Below we provide examples of the three levels of objection in the order of increasing specificity, starting with general objections to AI use per se, following with general objections to the use of specific AI tools, and concluding with objections to specific types of AIGC.
AIGC As Incompatible With Community Purpose
Ten participants reported opposing AIGC because they felt that using AI went against the purpose of their communities, which was to share posts created by humans. These participants objected to AI use in general, which suggests that there is no type of AIGC that they would welcome in their communities. These moderator’s opposition was ideological in nature, with one participant, a moderator of r/WritingPrompts, going so far as to say that they believed “[AIGC] is antithetical to everything that [our] subreddit is about”. These participants didn’t share objections to the quality of AIGC, but rather to the process through which it is created:
@@ 4.1.2. Opposition To AIGC
This moderator suggests that it is the personal, human-written nature of content in their community that differentiates them from other information sources like Google and Wikipedia. This quote suggests that the participant believes that contributors cannot “write in their own words” if they use AI, which we see as a context-specific norm that will likely differ between communities.
Current Generative AI Models Violate Community Values
Seven participants voiced ideological objections not to AI use per se, but rather to the use of current generative AI models, and by extension the content they produce, because they felt that certain characteristics of these models were incompatible with their communities’ values. While these characteristics of generative AI models may be difficult or unlikely to change, they are not inherent to the idea of AI, which suggests that these participants might not oppose AIGC in their communities were it produced by future generations of AI tools without these problematic characteristics. For example, one moderator of a creative subreddit shared that they object on ethical principles to AIGC because of the way that generative AI models are trained:
@@
This isn’t an objection to the use of AI per se, but rather to the use of AI tools that are trained according to the current paradigm, which the moderator feels is exploitative. This suggests that if model training methods evolve and address this concern then the moderator may no longer oppose AIGC in their community.
Another example is moderators’ stated opposition to generative AI models’ inability to reason, a trait which led a moderator of r/explainlikeimfive to describe ChatGPT as a “bullshit generator”. This is another objection based on process and not the content itself: even in instances where AIGC happens to be accurate, it is seen as problematic because it is produced by tools that work stochastically, as opposed to through reasoning. Some participants reported believing that generative AI models’ lack of relationship with the truth made them incompatible with the values of their communities:
Once again, this is not an objection to AI use per-se, but to the current paradigm of probabilistic generative AI models that are prone to hallucination. At this moment there may be no AI alternatives that would satisfy this objection, but this attitude does not close off future possibilities.
AIGC As Low Quality Content
Some participants didn’t object to AIGC in their communities on ideological grounds, but instead took a more practical stance and objected to AIGC because they saw it as low-quality content. While these were general attitudes, this opposition was often related to specific threats that they perceived from AIGC, which we will explore in greater detail in 4.2.
We heard from nine moderators who reported that AIGC did not meet the quality standards of their subreddits, like a moderator of r/AskHistorians who shared that AIGC, “tries to meet the substance and depth of a typical post…however, there are frequent glaring errors in both style and content.” Other moderators put it more bluntly, like a moderator of r/CryptoCurrency who noted that, “[AIGC] is not enjoyable to read”, or a moderator of r/Eldenring who shared: “[AIGC] is content that provides no value/discussion at all, it’s low effort content basically.” These participants did not take an ideological stance against AIGC, but evaluated it practically and saw it to add little value to their communities.
Other participants made practical objections to AIGC in their communities because they observed it being used in bad faith. Five participants, who moderated the largest subreddits, objected to AIGC because they saw it as a technique to rapidly accumulate Reddit karma [^6] https://support.reddithelp.com/hc/en-us/articles/204511829-What-is-karma, a practice known as karma-farming. Karma is a reputation points system that some communities use to limit who can make posts and karma-farming allows nefarious users to bypass this system. Some participants said they opposed AIGC because they had seen it used in the context of trolling, for example a moderator of r/politics shared that when they saw AIGC it was, “usually some form of trolling, or ‘outsourcing’ writing a comment the user wanted to make.” We also heard from participants who reported thinking that AIGC was primarily used to try and deceive other community members:

### v5

@@ 3. Methods: Interview Study
We chose to conduct semi-structured interviews with Reddit moderators, approaching the research questions qualitatively, to deeply “explore and understand the meaning individuals or groups ascribe to a social or human problem” (creswellResearchDesignQualitative2017a), specifically the newly introduced challenges of AIGC. We identified target communities by looking at their public stances towards AIGC, and recruited participants via outreach and snowball methods.
@@ 3.1. Recruitment
As our research goal was to explore moderator concerns, we sought to speak with moderators of communities who had restricted the use of AIGC. To identify such communities, we turned to public data about subreddits’ community rules. We performed a crawl of the Reddit website in the summer of 2023 and extracted the text-based rules from all subreddits indexed by Reddit’s “Top Communities” list [^1] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We crawled $n=337,399$ subreddits and found $n=87,596$ with non-empty rules and English as a primary language. We then performed a text search on the rules to identify subreddits with rules related to AIGC. While this method may be imprecise, it was sufficient for our purpose, which was to understand broad patterns to guide our recruitment strategy.
In the data that we gathered, 4% of all subreddits had rules about AI, but these rules were much more common in larger subreddits: when we restricted our analysis to top 1% of subreddits according to subscriber count, 20% had rules governing the use of AI. We thus designed a recruitment strategy focused on the largest subreddits, which appear to be the most concerned with the potential impact of AIGC. Many moderators of larger subreddits also moderate small forums and were able to speak to the differences between these types of communities.
We recruited and interviewed participants during the summer of 2023. We manually examined the rules of subreddits in descending order of subscriber count. When we found a subreddit with either an explicit rule governing the use of AI or a rule that might implicitly cover AI use we added it to our sample. We also looked for moderator discussions about AIGC in other public forums, such as the r/ModSupport subreddit, and added those moderators’ subreddits to our sample. We then messaged the moderators of each subreddit in our sample via Reddit’s Modmail [^2] https://support.reddithelp.com/hc/en-us/articles/210896606-What-is-Modmail- feature, sending Modmail recruitment messages to 60 subreddits. We used two additional techniques to recruit participants beyond the largest subreddits. First, we conducted snowball sampling and asked moderators to forward our recruitment message to other potential participants. Second, we posted our recruitment call on several social media platforms. To ensure respondents were actual moderators, we asked for their Reddit username and the subreddit that they moderate, then verified that the users were indeed listed as moderators via their public Reddit user profile. Participants were offered \$20 gift-cards as compensation. We conducted interviews on a rolling basis and continued recruiting until we had enough rich and complex data to adequately address our research questions, as per Braun and Clarke’s saturation criteria for Reflexive Thematic Analysis (braunSaturateNotSaturate2021). Overall, we interviewed fifteen participants, meeting or even exceeding the sample size norms of the HCI research community (caine16). These participants collectively moderated over 100 different subreddits with sizes ranging from less than ten members to more than 32 million (see participant information in Table 1).
@@ 3.2. Interviews
We designed an initial interview protocol based on our research questions and updated it as themes emerged during early interviews. Since pseudonymity is an important feature of the Reddit platform, we followed best practices in ethical online communities research (Fiesler24) and took special considerations to protect the privacy of our participants. We did not ask for or record participants’ names, demographic information, or any offline identifiers. As text is the primary mode of interaction on Reddit, we gave participants the opportunity to interview via text-based mediums, which is a common practice in Reddit and Social Computing research (Dimond12; Gilbert20; Jhaver19). Accordingly, participants were asked to choose a medium that was comfortable for them for a 60-minute, synchronous interview: video call, audio call, email exchange, or message exchange via the Reddit platform. Regardless of channel, we posed our research questions one at a time, in real time, and followed up in a semi-structured way. For example, for text-based interviews, we were online at the same time as our participants, exchanging messages in a back-and-forth manner, which allowed us to probe participants’ responses as necessary for deeper insight or clarity. The study was approved by our institution’s IRB. All participants read and accepted a consent agreement permitting their responses to be reproduced in research reports along with the participants’ associated subreddits. To ensure that we reliably represented participants’ views, we contacted participants for feedback and approval on their included quotes before publication, following CSCW community recommendations (mcdonald19). We were able to integrate all participants’ feedback in the final manuscript.
The first author conducted the interviews using these key guiding questions:
@@
    What concerns do moderators have about AIGC and how do they compare to other moderation concerns?
@@
For interviews conducted via audio and video calls, interview transcripts were generated automatically from recordings. The transcripts were manually reviewed and corrected by the first author. For the four interviews conducted via email and the four conducted via Reddit message, the text was captured in similarly formatted transcript files for analysis. Interviews lasted between 26 minutes and 147 minutes with a mean and median of 71 and 62 minutes, respectively. While we scheduled $60$ minutes for the interviews, some oral interviews ended early and some text-based interviews went longer, as participants took their time to respond to our questions: oral interviews had a mean duration of $47$ minutes ($SD=15$), compared to a mean duration of $91$ minutes ($SD=38$) for written interviews. Despite the semi-synchronous nature of the longer written interviews, all yielded meaningful conversations with at least eight back-and-forth question and answer exchanges. Written interviews served as a valuable complement to our oral conversations by providing rich data that was not shared via oral conversation, such as links to example AIGC posts from participants’ communities. Despite the differences between the mediums, both yielded deep and meaningful data that combine to create a more nuanced picture than any single medium could provide.
The first two authors then collaborated to iteratively code the interview content into categories, using an inductive thematic analysis method akin to Braun and Clarke’s Reflexive Thematic Analysis (braun21; Braun19). The two authors began by discussing themes that the first author observed while conducting the interviews and correcting the transcripts. From this conversation, an initial high-level set of codes emerged. Next, the first author open-coded all of the interview transcripts at the sentence-level, applying the initial codes and creating new ones where existing codes were insufficient. The two authors then reviewed the set of codes and refined them by merging similar codes and nesting those with a clear hierarchy. This round of coding did not produce new themes, only new sub-codes, suggesting that the themes were fairly stable. Both authors then independently applied these codes to the same three interview transcripts and compared their results. Neither application had created any new codes and the few cases of disagreement were resolved by further merging similar codes. This process yielded approximately 400 tagged phrases across all transcripts, which were grouped into the three high-level themes that we discuss below.
@@
1 r/todayilearned (32M) 25 Audio Call Amplified Attacks High
2 r/food (23M) 25 Reddit Message Exchange Lowered Quality High
3 r/explainlikeimfive (23M) 3 Email Exchange Less Accurate Information High
4 r/explainlikeimfive (22M) 2 Audio Call Less Accurate Information High
5 r/WritingPrompts (17M) 3 Video Call Fewer Opportunities for Skill Development High
6 r/politics (8.3M) 25 Reddit Message Exchange Strained Community Relationships Medium
7 r/CryptoCurrency (6.7M) 8 Reddit Message Exchange Amplified Attacks Medium
8 r/itookapicture (5.1M) 2 Email Exchange Fewer Opportunities for Skill Development Medium
9 r/NintendoSwitch (5.0M) 23 Email Exchange Lowered Quality Medium
10 r/Fantasy (3.4M) 1 Email Exchange Lowered Quality Medium
11 r/changemyview (3.4M) 2 Video Call Less Human Connection Low
12 r/changemyview (3.4M) 2 Audio Call Less Human Connection Low
13 r/AskHistorians (1.8M) 4 Video Call Less Accurate Information Medium
14 r/AskHistorians (1.8M) 1 Reddit Message Exchange Less Accurate Information Medium
15 r/GCTrading (22.7K) 3 Audio Call Amplified Attacks Low
@@ 4. Findings
Three main themes emerged from our interviews. With regards to RQ1, our interviews surfaced insights into moderators’ (1) concerns about AIGC in their communities; answering RQ2, our interviews elicited details on (2) how moderators’ communities are responding to AIGC and (3) the challenges of enforcing AIGC rules.
Our interviews did not reveal a single definition of “AIGC”. Instead, our participants mentioned three different categories of content related to our inquiry: images produced by generative AI, text produced (entirely or partially) by generative AI, and posts made by automated accounts (bots). Communities sometimes banned one of these categories while allowing others.
Participants generally used the term “AI-generated” in a pejorative way, but some acknowledged that AI-use falls along a spectrum, from benign to problematic, and shared the subjective factors that they considered when judging AI use. Some participants spoke of the intention of a poster [^3] We use the term poster to refer to both those posting top-level subreddit submissions as well as those posting comments in response to submissions., like a moderator of r/itookapicture [^4] To provide additional context, we introduce quotes with the name of the subreddit that the participant moderates. In the case where the participant moderates several subreddits, we include either the subreddit that the quote is referring to or, for more general quotes, the largest subreddit that the participant moderates. who said: “I don’t mind people using AI to better state their thoughts nor create art as long as it is not being used deceptively.” Others referenced the degree to which a poster relied on AI, like a moderator of r/ChangeMyView who reported being OK with a post made with AI, “if it was sufficiently [a poster’s own] words, just sort of juiced up by ChatGPT.” We heard something similar from a participant who cared that a post was a poster’s “own intellectual contribution or content”:
@@
Other participants took a less nuanced view, like a moderator of r/WritingPrompts, who relayed their community’s official stance: “Let’s be absolutely clear: you are not allowed to use AI in this subreddit, you will be banned”. Despite hearing arguments that, “ChatGPT is another writing tool that authors are using and banning ChatGPT will be the same as banning a spellchecker,” this participant disagreed and categorically rejected all uses of AI, going so far as to say that, “[AIGC] is antithetical to everything that [our] subreddit is about.” The simplicity of this stance has its appeal, as more subjective definitions of AIGC introduce the possibility that community members will disagree about specific examples—an implication which we will revisit in later sections.
4.1. Findings: Moderator Concerns About AIGC (RQ1)
Our recruitment focused on moderators of communities that are actively restricting the use of AIGC. Accordingly, while our participants offered differing definitions of AIGC, they generally opposed its use in their communities. We group our participants’ reported AIGC concerns into three, often overlapping, categories: content quality concerns, social dynamics concerns, and governance concerns. Despite general opposition, participants also acknowledged some potential benefits of AIGC, detailed below.
Table 2. Summary of our participants’ main concerns about AIGC in their communities.
Concern Examples
Low Quality Content Poor prose, factual inaccuracy, off topic
Impacted Social Dynamics Fewer authentic connections, strained relationships, ideological opposition
Difficult to Govern Increased volume of community attacks, harder to detect deception
Table 2 summarizes the three main categories of concerns that our participants shared. Content quality concerns take issue with the quality of AIGC and worry that it decreases the utility of an online community. Social dynamics concerns focus on how AIGC reduces the social value that an online community provides its members. Governance concerns emphasize the ways in which AIGC makes the job of moderating online communities more difficult. Naturally, these categories overlap and interact, and our participants often described considerations that span multiple categories.
Moderators shared various perspectives on the magnitude of these concerns. When asked if AIGC was one of their top concerns, a moderator for r/CryptoCurrency noted that, “Early 2023, I would’ve said yes, but right now it has calmed down quite a bit.” On the other hand, some participants did regard AIGC as a major concern:
r/news moderator: The concern is the more weaponized stuff…the folks who are utilizing these tools for very focused purposes, often for political reasons, or for commercial reasons, those are the ones that I think concern us the most. Whether or not somebody manages to get a post up there that doesn’t belong for a few hours, most of us don’t care…But what we do care about is creating a veneer of legitimacy for bad actors.
We will explore the challenges of detecting AIGC in Section 4.3, but first we dig deeper into the types of concerns that our participants shared.
4.1.1. Content Quality Concerns
First and foremost, moderators were concerned about the quality of AIGC. Nine participants reported that AIGC did not meet the quality standards of their subreddits, like a moderator of r/AskHistorians who said that AIGC, “tries to meet the substance and depth of a typical post…however, there are frequent glaring errors in both style and content.” Participants shared several reasons that they considered AIGC low quality, such as their issues with its style, its tendency to be factually inaccurate, and the belief that it was off-topic.
Several participants reported that there were certain stylistic issues with AIGC that made it low-quality content. We heard from one moderator that AI-generated answers to questions in their community often, “do not address the question being asked…tend to be very general and ‘hedge’ more than a real human.” Other moderators put it more bluntly, like a moderator of r/CryptoCurrency who noted that, “[AIGC] is not enjoyable to read”, or a moderator of r/Eldenring who told us: “[AIGC] is content that provides no value/discussion at all, it’s low effort content basically.” This moderator considered effort to be a signal of value or quality, which we will revisit in the discussion. They went on to share that this sort of low effort content can, “[discourage] users who want to post their own original content”.
Four participants who moderate communities that emphasize the accuracy of information (e.g. r/AskHistorians and r/explainlikeimfive) specifically mentioned their concern about inaccurate AI hallucinations. This concern is captured by a moderator of r/explainlikeimfive who objected to AIGC on the grounds that ChatGPT was a “bullshit generator”, or by another participant who shared that AIGC’s inaccuracy was unacceptable in their communities:
This participant described AIGC as “truthy-seeming” because of its tendency to include fake citations and felt that the appearance of being thoroughly researched increased the chances that readers would believe it, which, they feared, would “undermine the trust that people have [in the community].”
Finally, we heard from several moderators that AIGC was simply off-topic in their communities. For example, a moderator of r/itookapicture shared that the goal of their community was, “sharing great photography and photography techniques,” and accordingly, “preventing the submission of AI generated images and human-created compositions alike are a top priority.”
4.1.2. Social Dynamics Concerns
Of course, online communities are not only a source of high-quality content, but also of meaningful social interactions. We heard from ten participants who were concerned that AIGC would negatively impact the social dynamics in their communities, reducing the social value that their communities provide to members. Participants mentioned several ways that this could happen, such as by decreasing opportunities for human connection, straining relationships, and violating their communities’ values.
Ten participants reported that AI went against their communities’ purpose, which was to share posts created by humans. These participants mentioned the inherently inhuman nature of AIGC and expressed concern that this would lead to less authentic connection in their communities. These were not necessarily objections to the quality of AIGC, but rather to the process through which it is created:
@@ 4.1.2. Social Dynamics Concerns
These participants believed people come to their community to connect with others through dialogue and that AIGC does not meaningfully contribute to—or even detracts from—such conversations. A moderator from r/changemyview summarized this sentiment well: “How can we change your view when you haven’t actually stated your view at all? This is ChatGPT’s view. And if I wanted to, I could go argue with ChatGPT. But that’s not helping here.” We heard something similar from six moderators of communities of practice (e.g r/itookapicture, which is devoted to photography) who felt AIGC did not contribute to the collaborative learning that the community aimed to encourage. These are communities where members seek to develop particular skills by doing, sharing, and receiving feedback from other practitioners. Accordingly, posts produced without using their practice did not provide educational value to the community. This concern was summarized by a moderator of r/WritingPrompts who shared that their community existed to, “[give] writers a chance to practice their craft and practice their skill. And if you are taking away the practice element of that, because all you’re doing is feeding a prompt into ChatGPT…The server is no longer serving its purpose.”
Additionally, we heard concern from participants that AIGC could strain relationships in their community. For example, a moderator of a political subreddit noted that: “the possibility [of AIGC] fuels Redditors accusing other comments of being written by AI (which is a form of incivility).” Beyond incivility, the controversial nature of AIGC could cause discussions to devolve into off-topic arguments:
Finally, some participants took an ideological stance and objected to the use of AI and, by extension, AIGC, because they saw it as incompatible with their communities’ values. For example, we heard from a moderator of a creative subreddit who objected specifically to the way that generative AI models are trained:
@@
4.1.3. Governance Concerns
Participants also objected to AIGC because it made the job of moderating their communities more difficult. We heard from eight participants who saw AIGC being used for malicious purposes and were concerned that existing moderation techniques would struggle to handle it. These moderators spoke of pre-existing problems, such as spam and harassment, that are more difficult to manage when they are amplified in scale or sophistication by AIGC. While AIGC may be less common than other moderation challenges, it can still be quite disruptive:
Five participants, who moderated the largest subreddits, spoke of AIGC being used in broad attacks on their communities. These attacks often involved automated accounts posting large amounts of spam. These attacks may aim to disrupt the functioning of a community through sheer volume, as we heard from one moderator who experienced a ChatGPT-powered spam attack that required the help of Reddit staff to resolve:
r/AskHistorians moderator: A few months ago, we had an instance where we were subject to a bot attack that was using ChatGPT…we ended up reporting it to the admins who were able to take care of it through whatever they do on their back end.
Other times these attacks aim to deceive users as a part of influence operations, as we heard from a moderator of r/news, one of the platform’s largest subreddits, who shared their experience dealing with, “a highly sophisticated deployment, an attempt to sway a narrative by utilizing accounts that had had their karma bolstered by automated commenting.” This participant references Reddit karma [^6] https://support.reddithelp.com/hc/en-us/articles/204511829-What-is-karma, a reputation points system that communities can use to limit who can post, noting that AIGC can be used to bypass this system through a practice known as karma-farming, in which automated accounts make frequent posts in order to rapidly accumulate karma. Karma-farming is not a new practice, but this moderator thought that AIGC made it harder to detect, and thus more effective. Beyond karma-farming, we also heard about AIGC being used in other forms of deception: